let k be an integer and p is a prime such that the quadratic equation x^2+kx+p=0 has two distinct positive integer solutions. The value of (k+p)=,,?
Answer by ankor@dixie-net.com(22740)

let k be an integer and p is a prime such that the quadratic equation
x^2 + kx + p = 0 has two distinct positive integer solutions.
The value of (k+p)=,,?
:
If p is prime and it has two integer solutions, one of the solutions has to be 1
therefore k = p + 1
x^2 + (p+1)x + p = 0
" The value of (k+p)="
(p+1) + p = 2p+1